If a business borrows $15,000 and repays $21,300 in 4 years, what is the simple interest rate?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

I = P*R*N
where P= principal amount=15000
r= rate of interest = r
n= number of years = 4
Interest = Amount - Principal
Interest = 21300-15000=6300
6300 = 15000*4*r
6300/(15000*4)=r
r= 0.105
in percent it will be 10.5

Answer by stanbon(75887)

If a business borrows $15,000 and repays $21,300 in 4 years, what is the simple interest rate?
--------
A(t) = P(1+rt)
21300 = 15000(1 + r*4)
1.42 = 1 + 4r
4r = 0.42
r = 0.1050 = 10.5%
======================
Cheers,
Stan H.
================